OpenStudy (anonymous):

I'm stuck on this question, can anyone help? Find the equation of a straight line graph passing through (-5,10) and (5,20)

OpenStudy (anonymous):

in the form of y=mx+b?

OpenStudy (paounn):

There are several options. 1: \[{{x-(-5)} \over { 5 - (-5)} } = {{y- 10} \over {20-10}} \] it takes some basic calculation but it gives the answer. 2:\[\det \left[\begin{matrix}x & y & 1 \\ (-5) & 10 & 1\\ 5& 10 &1 \end{matrix}\right] = 0\] if you know some linear algebra. 3: pick a generic line y= mx+q. If the first point is on the line, then 10 = m (-5) +q. If the second point belongs to it as well, it must be 20 = m (5) +q. You have a linear system in m and q. Pick your poison